# Boosted Neural Decoders: Achieving Extreme Reliability of LDPC Codes for 6G Networks

### v1

@@ III-A Boosting Learning Using Uncorrected Vectors
For the training of the base decoder, we follow the conventional training method [9]: the base decoder is trained with received vectors sampled at the waterfall region, ${\rm E}_{\rm b}/{\rm N}_{0}$ values $\{2.0,2.5,3.0,3.5,4.0\}$ (dB). Also, we employ the spatial weight sharing (i.e., ${\bf w}_{\text{B}}=\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell=1}^{\ell_{1}}$) because it shows a similar performance with the full weight set in the waterfall region. The weights are initially set to $1$, which means that the NMS decoder is equivalent to the MS decoder before training.
Fig. 4(a) compares the decoding performance with ¯$\overline{\ell}=20$ for the base NMS, MS, and WMS decoders. The WMS decoder uses a single weight of $0.75$. While the WMS decoder shows superior waterfall performance to the MS decoder, it exhibits an severe error floor. The base NMS decoder, trained in the waterfall region, shows waterfall performance that is similar to that of the WMS decoder. However, its performance deteriorates compared to the MS decoder in the error floor region. To improve the error floor performance, one can extend the training region to include the error floor region. However, this approach is ineffective because of the almost zero FER loss in the error floor region. For example, the FER of the MS decoder (initial state of the NMS decoder) at $4.5$ dB is around $2\times 10^{-5}$, which implies that only a small fraction of training samples produce a non-zero loss.
To effectively train the post decoder, we accumulate training samples by exclusively gathering UC vectors that the base decoder fails to correct. A UC vector consists of channel LLRs leading to a decoding failure. Subsequently, the post decoder is trained with the UC vectors. In terms of neural networks, the network is constructed up to $\overline{\ell}$, and the weights for iterations to $\ell_{1}$ are fixed by the weight set ${\bf w}_{\text{B}}$ and the weight set ${\bf w}_{\text{P}}$ for iterations from $\ell_{1}+1$ to $\overline{\ell}$ is trained. For this experiment, we utilize the full weight set ${\bf w}_{\text{P}}=\{\overline{w}^{(\ell)}_{v_{p}},w^{(\ell)}_{c_{p}\rightarrow v_{p}}\}_{\ell=\ell_{1}+1}^{\overline{\ell}}$ to observe the best performance, and we will apply weight sharing to ${\bf w}_{\text{P}}$ in the following subsection.
@@
For Cases 1 and 2, we collect $60{,}000$ UC vectors, allocating $50{,}000$ for training, $5{,}000$ for validation, and $5{,}000$ for testing. Fig. 4(b) presents the distribution of the number of bit errors in decoded vectors after base and post decoding for the test samples. Observing Case 1 after base decoding, the majority of errors are concentrated in a small count, implying that most errors in the error floor region arise from small trapping sets. For convenience, we refer to vectors with $11$ or fewer errors as small-error vectors. In Case 1 after base+post decoding, it is evident that the post decoder successfully corrects most of the small-error vectors. The post decoder exhibits an FER of $0.322$ among the $5{,}000$ test samples, denoted as the test FER. This results in a reduction of the FER from the base decoder to the cascaded decoder (base+post) by a factor of $0.322$ at $4.5$ dB, as shown in Fig. 4(a). The test FER will serve as an important performance metric throughout this paper.
TABLE II: Test BERs and FERs for representative loss functions.
@@
In contrast, Case 2 after base decoding shows a widespread distribution of error counts, as it collects UC vectors at the waterfall region. This distribution remains largely unchanged even after post decoding. Therefore, the strategy for Case 1 is more effective in reducing the error floor, as clearly shown in Fig. 4(a). In other words, we can conclude that collecting training samples at the error floor region is vital for training decoders specialized in combating the error floor.
@@
Before proceeding to the next subsection, we compare the loss functions in Table II, assuming the boosting learning with Case 1. The result shows that the FER loss function, as expected, exhibits superior test FER performance although the BER performance is worse than that of the soft BER loss function. Since the problem of the error floor region is the gradual decrease in the FER curve rather than the BER, we choose to utilize the FER loss function.
@@ III-B Block-wise Training Schedule
Figure 5: (a) illustration of the block-wise training schedule and (b) comparison of test FERs across iterations.
TABLE III: Test FERs of the block-wise training with various $\Delta_{1}$ and $\Delta_{2}$.
@@
To further enhance error floor performance, we consider a larger number of iterations, $(\ell_{2}=30,\overline{\ell}=50)$, compared to $(\ell_{2}=10,\overline{\ell}=30)$ used in the previous subsection. However, this induces deep hidden layers which, in turn, exacerbates the vanishing gradient problem. To address this problem, we propose a block-wise training schedule as illustrated in Fig. 5(a). The training schedule incrementally trains weights associated with $\Delta_{1}$ iterations per stage. Initially, weights for the first $\Delta_{1}$ iterations are trained, and then the focus shifts to the next $\Delta_{1}$ iterations. The weights of previous $\Delta_{2}$ iterations are retrained using the trained weights in earlier stage as initial weights. This retraining is introduced to help the training process avoid local minima. Note that the one-shot training [9] belongs to a special case of $\Delta_{1}=\ell_{2},\Delta_{2}=0$, and the iter-by-iter method [20] is of $\Delta_{1}=1$, $\Delta_{2}=0$.
To optimize the block-wise training parameters $\Delta_{1}$ and $\Delta_{2}$, we evaluate the test FER for varying configurations of these parameters as shown in Table III. For scenarios without retraining (i.e., $\Delta_{2}=0$), increasing $\Delta_{1}$ from $1$ to $10$ achieves performance gains. This can be attributed to the training of more weights simultaneously, aiding in escaping from local minima. Nonetheless, excessively large values of $\Delta_{1}$ can result in undertraining of the weights due to the vanishing gradient problem, thereby degrading performance. Table III indicates that the optimal result is obtained with $\Delta_{1}=5$ and $\Delta_{2}=10$.
Fig. 5(b) shows the evolution of test FERs according to iteration numbers. For the one-shot training [9], the vanishing gradient problem impedes the effective training of early iteration weights. This results in a consistent test FER up to iteration $40$ for ¯$\overline{\ell}=50$. Since this behavior is observed for ¯$\overline{\ell}=40$ as well, the test FERs for ¯$\overline{\ell}=40$ and ¯$\overline{\ell}=50$ are similar. The iter-by-iter method [20] optimizes weights for each iteration one-by-one, reducing the test FER from the initial iteration. However, this approach gets trapped in a local minimum and leads to only slow FER improvements over iterations. A similar trend is observed with the multi-loss method [9].
@@ III-C Dynamic Weight Sharing
TABLE IV: Test FERs for various weight sharing techniques.
@@
Fig. 6 shows the trained weights using our sharing technique. With only three distinct weights per iteration – VN weight (VW) $\overline{w}^{(\ell)}$, CN weight (CW) $w^{(\ell)}$, and UCN weight (UCW) $\hat{w}^{(\ell)}$– the weight evolution can be depicted simply in a two-dimensional graph. In the base stage, the CW and VW values are roughly around $0.75$ and $1$, respectively, which are the weights of the WMS decoder. As a result, the WMS and NMS decoders perform similarly in the waterfall region. On the contrary, a significant change in weights is observed in the post stage. At iteration $21$, the VW increases and CW decreases substantially. This means that the channel LLR values are given more weight when performing the sum operation (1), while the messages coming from CNs are attenuated. From iteration $22$ onward, the CW gradually increases and the VW decreases, suggesting an increasing dependence on CN messages. This dynamic weight evolution enhances decoding diversity in the post decoding stage that lowers the error floor.
The method differing UCW and SCW has been explored in prior works [38, 46]. However, the method in [46] employs a single suppression factor $\rho$ to modify the UCW based on the CW (i.e., ${\hat{w}}^{(\ell)}=(1+\rho)w^{(\ell)}$), limiting the capability to manage various error patterns effectively. Meanwhile, the scheme in [38] assigns the UCN weight to all $z$ CNs derived from a single proto CN if any one of them is unsatisfied. This is not suitable for correcting small-error vectors with few UCNs since it does not consider individual CNs distinctively, whereas we pinpoint specific UCNs and apply weights individually.
III-D Transfer Learning
@@
While conventional learning methods using received vectors as training samples have an immediate sampling process, the proposed boosting learning requires numerous decoding trials to collect UC vectors, leading to a time-intensive sampling process. For example, obtaining $T$ UC vectors requires $T$/$\rm FER$ decoding trials in average. In previous sections, we collect $50{,}000$ UC vectors at the base decoding FER of $2\times 10^{-5}$. This sampling process requires $2.5\times 10^{9}$ decoding trials on average. If we require an extremely low FER and training of multiple codes, the sampling process may become infeasible.
@@
Figure 8: Comparison of test FERs according to the number of training samples.
Fig. 8 shows the test FERs according to the number of training samples used for the post decoder with $15$ iterations and the WiMAX LDPC code. Without transfer learning and using the full weight set $\{\overline{w}_{v_{p}}^{(\ell)},w_{c_{p}\rightarrow v_{p}}\}_{\ell}$ (block dotted lines), the test FER begins at $0.56$ with zero training sample (i.e., no training) and reduces to $0.30$ after using $10{,}000$ samples. Though not shown in the figure, the test FER stabilizes at $0.21$ with $50{,}000$ samples. Using shared weights, $\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell}$ for spatial sharing and $\{\overline{w}^{(\ell)},w^{(\ell)},{\hat{w}}^{(\ell)}\}_{\ell}$ for dynamic sharing, the test FERs are saturated to their maximum performances with $10{,}000$ samples. This shows another advantage of weight sharing: fewer trainable weights lead to fewer required sample counts.
TABLE V: Test FERs for transfer learning with various variations
@@
For transfer learning, the initial weights for the WiMAX LDPC code of $M=6,N=24,E=88,R=3/4$ are set by the pre-trained weights for the WiFi LDPC code $M=4,N=24,E=88,R=5/6$. Since both the WiFi and WiMAX codes have the same number of edges $E$ and nodes $N$, a direct transfer is available without weight sharing. Nonetheless, as shown in Fig. 8, the performance of the full weight set converges at $10{,}000$ samples (black solid line), while using weight sharing significantly reduces the required number of samples to $1{,}000$ (blue and red solid lines). Notably, the initial test FERs with weight sharing are already low even before training. A small amount of fine-tuning is enough to obtain the optimal weight set. As a result, the combined use of transfer learning and weight sharing reduces the required number of samples from $50{,}000$ to $1{,}000$, thereby accelerating the sampling process by a factor of $50$.
Table V presents a range of experiments to show the effectiveness of transfer learning across different code types, code rates, channels, and decoders. The target conditions for transfer learning are the WiMAX LDPC code with code rate of $3/4$, the AWGN channel, and the 5-bit quantized decoder, which have been assumed as a running example so far. Among them, we modify one condition at a time and derive the pre-trained weights. For the code type, we use the pre-trained weights from the WiFi LDPC code, whose results are also in Fig. 8. For all cases in Table V, the fine tuning with $1{,}000$ samples improves the test FER compared to the baseline FER with merely transferring of weights, showing the effectiveness of transfer learning. In addition, this result demonstrates the flexibility of the proposed boosting learning method that can be applicable to diverse code, channel, and decoder types.
III-E Data Augmentation
TABLE VI: Comparison between the plain sampling method and the accelerated sampling method using transfer learning and data augmentation
@@ III-E Data Augmentation
$\{\overline{w}_{v_{p}}^{(\ell)},w_{c_{p}\rightarrow v_{p}}\}_{(\ell)}$ Non-transfer $50000$ $2\times 10^{-5}$ N/A N/A $2.5\times 10^{9}$ $1$ 0.209
@@
In this subsection, we introduce the use of importance sampling for data augmentation to further speed up the sampling process. Importance sampling was used in estimating error floor performance of LDPC codes [7, 15]. Under the assumption of the binary phase shift keying (BPSK) modulation and all-zero codeword, stronger noise to variable nodes (VNs) within the trapping set is introduced by subtracting $\beta$ from their received values $y_{i}$. (i.e., $y_{i}=1-\beta+n_{\rm ch}$ for $i$ within the trapping set, where $n_{\rm ch}$ is the channel noise.) This approach allows for the rapid evaluation of a relatively high FER, which is used to estimate the actual FER. Thus, we can estimate an extremely low actual FER with a small number of decoding trials.
@@
The proposed data augmentation method also generates intensified noise at specific bit positions. As depicted in Fig. 7, the data augmentation block processes $T$ input UC vectors to produce $T\times D$ augmented UC vectors, where $T=100$ and $T\times D=1{,}000$. For each input UC vector, the base decoding is performed to identify the positions of error bits. These error bit positions, denoted as $\mathcal{E}$, are identified at the iteration when the number of UCNs is minimal. Subsequently, we accumulate $D$ UC vectors using the base decoder under a modified channel model where received vectors are generated from the AWGN channel as usual, and received values for VNs in $\mathcal{E}$ are shifted by $-\beta$. This shift increases the probability of decoding failures and enables the rapid collection of $D$ additional UC vectors that share characteristics with the input UC vector. Each UC vector entering the data augmentation block is augmented into $D$ UC vectors, thereby generating a total of $T\times D$ UC vectors from $T$ input UC vectors.
@@
Table VI shows how transfer learning and data augmentation accelerates the sampling process. Without weight sharing, collecting $50{,}000$ UC vectors (indicated in ($a$)) with the FER of the base decoder (indicated in FER ($b$)) requires $a/b=2.5\times 10^{9}$ decoding trials using the plain sampling method. By applying the dynamic weight sharing technique $\{\overline{w}^{(\ell)},w^{(\ell)},{\hat{w}}^{(\ell)}\}_{\ell}$, the required number of UC vectors is reduced to $10{,}000$ without transfer learning and further to $1{,}000$ with the addition of transfer learning. For data augmentation, Table VI presents results for $T=100$, $D=10$, and various $\beta$ values. With data augmentation, the augmented UC vectors ($c$) equal to $T\times D=1{,}000$, and the FER ($d$) represents to the FER under the modified channel with intensified noise. The decoding trials in data augmentation are $c/d$. Thus, the total number of decoding trials is given by $a/b+c/d$. The weight sharing technique alone reduces the required trials by a factor of $5$ and accelerates the sampling speed by the same factor. Transfer learning adds a 10-fold increase in speed, and this gain is further increased by data augmentation. As $\beta$ increases, the FER ($d$) in data augmentation increases, giving more gains. However, accumulated samples with high $\beta$ might differ in characteristics from those collected through plain sampling, which leads to degraded test FER performance. Therefore, we opt for $\beta=0.7$, achieving a ×$\times 361$ increase in sampling speed without significant loss in the test FER.
@@
Additionally, to eliminate the need for validation samples, we adopt the step decaying learning rate strategy. Until now, $5{,}000$ UC vectors has been allocated for validation samples. However, since the NMS decoder is not prone to overfitting on training data, the weights optimized on validation samples have similar performance to those at the last epoch. To further enhance stability in convergence, we use the step decaying learning rate, starting at $0.001$ and reducing it by half every $20$ epochs, and select the weights at the last epoch without validation samples.

### v2

@@ III-A Boosting Learning Using Uncorrected Vectors
For the training of the base decoder, we follow the conventional training method [9]: the base decoder is trained with received vectors sampled at the waterfall region, ${\rm E}_{\rm b}/{\rm N}_{0}$ values $\{2.0,2.5,3.0,3.5,4.0\}$ (dB). Additionally, we employ the spatial weight sharing (i.e., ${\bf w}_{\text{B}}=\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell=1}^{\ell_{1}}$) because it shows a similar performance with the full weight set in the waterfall region. The weights are initially set to $1$, which means that the NMS decoder is equivalent to the MS decoder before training.
Fig. 4(a) compares the decoding performance with ¯$\overline{\ell}=20$ for the base NMS, MS, and WMS decoders. The WMS decoder uses a single weight of $0.75$. While the WMS decoder shows superior waterfall performance to the MS decoder, it exhibits a severe error floor. The base NMS decoder, trained in the waterfall region, shows waterfall performance that is similar to that of the WMS decoder. However, its performance deteriorates compared to the MS decoder in the error floor region. To improve the error floor performance, one can extend the training region to include the error floor region. However, this approach is ineffective because of the almost zero FER loss in the error floor region. For example, the FER of the MS decoder (initial state of the NMS decoder) at $4.5$ dB is around $2\times 10^{-5}$, which implies that only a small fraction of training samples produce a non-zero loss.
To effectively train the post decoder, we accumulate training samples by exclusively gathering UC vectors that the base decoder fails to correct. A UC vector consists of channel LLRs leading to a decoding failure. Subsequently, the post decoder is trained with the UC vectors. In terms of neural networks, the network is constructed up to $\overline{\ell}$, and the weights for iterations up to $\ell_{1}$ are fixed by the weight set ${\bf w}_{\text{B}}$ and the weight set ${\bf w}_{\text{P}}$ for iterations from $\ell_{1}+1$ to $\overline{\ell}$ is trained. For this experiment, we utilize the full weight set ${\bf w}_{\text{P}}=\{\overline{w}^{(\ell)}_{v_{p}},w^{(\ell)}_{c_{p}\rightarrow v_{p}}\}_{\ell=\ell_{1}+1}^{\overline{\ell}}$ to observe the best performance, and we will apply weight sharing to ${\bf w}_{\text{P}}$ in the following subsection.
@@
For Cases 1 and 2, we collect $60{,}000$ UC vectors, allocating $50{,}000$ for training, $5{,}000$ for validation, and $5{,}000$ for testing. Fig. 4(b) presents the distribution of the number of bit errors in decoded vectors after base and post decoding for the test samples. Observing Case 1 after base decoding, the majority of errors are concentrated in a small count, implying that most errors in the error floor region arise from small trapping sets. For convenience, we refer to vectors with $11$ or fewer errors as small-error vectors. In Case 1 after base+post decoding, it is evident that the post decoder successfully corrects most of the small-error vectors. The post decoder exhibits FER of $0.322$ among the $5{,}000$ test samples, denoted as test FER. This results in a reduction of FER from the base decoder to the cascaded decoder (base+post) by a factor of $0.322$ at $4.5$ dB, as shown in Fig. 4(a). The test FER will serve as an important performance metric throughout this paper.
TABLE II: Test BER and FER for representative loss functions.
@@
In contrast, Case 2 after base decoding shows a wider distribution of error counts, as it collects UC vectors at the waterfall region. This distribution remains largely unchanged even after post decoding. Therefore, the strategy for Case 1 is more effective in reducing the error floor, as clearly shown in Fig. 4(a). In other words, it can be concluded that collecting training samples from the error floor region is vital for training decoders specialized in combating the error floor. The post decoder is trained to introduce a different form of diversity compared to the base decoder, addressing the error floor caused by a small number of errors.
@@
Before proceeding to the next subsection, we compare the loss functions in Table II, assuming the boosting learning with Case 1. The result shows that the FER loss function, as expected, exhibits superior test FER performance although the BER performance is worse than that of the soft BER loss function. Since the problem of the error floor region is the gradual decrease in the FER curve rather than the BER curve, we choose to utilize the FER loss function.
@@ III-B Block-wise Training Schedule
Figure 5: (a) Illustration of the block-wise training schedule and (b) comparison of the test FER across iterations.
TABLE III: Test FER of the block-wise training with various $\Delta_{1}$ and $\Delta_{2}$.
@@
To further enhance error floor performance, we consider a larger number of iterations, $(\ell_{2}=30,\overline{\ell}=50)$, compared to $(\ell_{2}=10,\overline{\ell}=30)$ used in the previous subsection. However, this induces deep hidden layers which, in turn, exacerbates the vanishing gradient problem. To address this problem, we propose a block-wise training schedule as illustrated in Fig. 5(a). The training schedule incrementally trains weights associated with $\Delta_{1}$ iterations per stage. Initially, weights for the first $\Delta_{1}$ iterations are trained, and then the focus shifts to the next $\Delta_{1}$ iterations. The weights of previous $\Delta_{2}$ iterations are retrained using the trained weights from earlier stages as initial weights. This retraining is introduced to help the training process avoid local minima. Note that the one-shot training [9] belongs to a special case of $\Delta_{1}=\ell_{2},\Delta_{2}=0$, and the iter-by-iter method [20] is of $\Delta_{1}=1$, $\Delta_{2}=0$.
To optimize the block-wise training parameters $\Delta_{1}$ and $\Delta_{2}$, we evaluate test FER for varying configurations of these parameters as shown in Table III. For scenarios without retraining (i.e., $\Delta_{2}=0$), increasing $\Delta_{1}$ from $1$ to $10$ achieves performance gains. This can be attributed to the training of more weights simultaneously, aiding in escaping local minima. Nonetheless, excessively large values of $\Delta_{1}$ can result in undertraining of the weights due to the vanishing gradient problem, thereby degrading performance. Table III indicates that the optimal result is obtained with $\Delta_{1}=5$ and $\Delta_{2}=10$.
Fig. 5(b) shows the evolution of test FER according to iteration numbers. For the one-shot training [9], the vanishing gradient problem impedes the effective training of early iteration weights. This results in consistent test FER up to iteration $40$ for ¯$\overline{\ell}=50$. Since this behavior is observed for ¯$\overline{\ell}=40$ as well, the test FER values for ¯$\overline{\ell}=40$ and ¯$\overline{\ell}=50$ are similar. The iter-by-iter method [20] optimizes weights for each iteration one-by-one, reducing test FER from the initial iteration. However, this approach gets trapped in a local minimum and leads to only slow FER improvements over iterations. A similar trend is observed with the multi-loss method [9].
@@ III-C Dynamic Weight Sharing
TABLE IV: Test FER for various weight sharing techniques.
@@
Fig. 6 shows the trained weights using our sharing technique. With only three distinct weights per iteration – ChW $\overline{w}^{(\ell)}$, CW $w^{(\ell)}$, and UCW $\hat{w}^{(\ell)}$– the weight evolution can be depicted simply in a two-dimensional graph. In the base stage, the CW and ChW values are roughly around $0.75$ and $1$, respectively, which are the weights of the WMS decoder. As a result, the WMS and NMS decoders perform similarly in the waterfall region. On the contrary, a significant change in weights is observed in the post stage. At iteration $21$, the ChW increases and CW decreases substantially. This means that the channel LLR values are given more weight when performing the sum operation (1), while the messages coming from CNs are attenuated. From iteration $22$ onward, the CW gradually increases and the ChW decreases, suggesting an increasing dependence on CN messages. This dynamic weight evolution enhances decoding diversity in the post decoding stage that lowers the error floor.
The method differing UCW and CW has been explored in prior works [46, 55]. However, the method in [55] employs a single suppression factor $\rho$ to modify the UCW based on the CW (i.e., ${\hat{w}}^{(\ell)}=(1+\rho)w^{(\ell)}$), limiting the capability to manage various error patterns effectively. Meanwhile, the scheme in [46] assigns the UCW to all $z$ CNs derived from a single proto CN if any one of them is unsatisfied. This is not suitable for correcting small-error vectors with few UCNs since it does not consider individual CNs distinctively, whereas we pinpoint specific UCNs and apply weights individually.
III-D Ablation Study
TABLE V: Ablation study for the $(n=576,k=432)$ WiMAX LDPC code.
Learning Schedule Sharing
O: Boosting O: Block-wise O: Dynamic FER
X: Conv. X: One-shot X: No sharing (${\rm E}_{\rm b}/{\rm N}_{0}$ 5.0dB)
X X X $3.14\times 10^{-6}$
O X X $2.77\times 10^{-7}$
O O X $1.84\times 10^{-7}$
O O O $1.85\times 10^{-7}$
The ablation study for the techniques introduced so far is presented in Table V. For the $(n=576,k=432)$ WiMAX LDPC code, using boosting learning lowers FER at ${\rm E}_{\rm b}/{\rm N}_{0}=5.0$ dB by more than one order of magnitude, and block-wise training reduces it by approximately $1.5$ times. Additionally, dynamic sharing significantly reduces the number of weight sets with only minimal degradation in performance. In summary, the most essential technique for improving performance among the proposed methods is boosting learning.
III-E Transfer Learning
@@
While conventional learning methods using received vectors as training samples have an immediate sampling process, the proposed boosting learning requires numerous decoding trials to collect UC vectors, leading to a time-intensive sampling process. For example, obtaining $T$ UC vectors requires $T$/$\rm FER$ decoding trials on average. In previous sections, we collect $50{,}000$ UC vectors at the base decoding FER of $2\times 10^{-5}$. This sampling process requires $2.5\times 10^{9}$ decoding trials on average. If we require extremely low FER and training of multiple codes, the sampling process may become infeasible.
@@
Figure 8: Comparison of the test FER according to the number of training samples.
Fig. 8 shows the test FER according to the number of training samples used for the post decoder with $15$ iterations and the WiMAX LDPC code. Without transfer learning and using the full weight set $\{\overline{w}_{v_{p}}^{(\ell)},w_{c_{p}\rightarrow v_{p}}^{(\ell)}\}_{\ell}$ (black dotted lines), the test FER begins at $0.56$ with zero training sample (i.e., no training) and reduces to $0.30$ after using $10{,}000$ samples. Though not shown in the figure, the test FER stabilizes at $0.21$ with $50{,}000$ samples. Using shared weights, $\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell}$ for spatial sharing and $\{\overline{w}^{(\ell)},w^{(\ell)},{\hat{w}}^{(\ell)}\}_{\ell}$ for dynamic sharing, the test FER is saturated at their maximum performances with $10{,}000$ samples. This shows another advantage of weight sharing: Fewer trainable weights lead to fewer required sample counts.
TABLE VI: Test FER for transfer learning with various variations
@@
For transfer learning, the initial weights for the WiMAX LDPC code of $M=6,N=24,E=88,R=3/4$ are set by the pre-trained weights for the WiFi LDPC code $M=4,N=24,E=88,R=5/6$. Since both the WiFi and WiMAX codes have the same number of edges $E$ and nodes $N$, a direct transfer is available without weight sharing. Nonetheless, as shown in Fig. 8, the performance of the full weight set converges at $10{,}000$ samples (black solid line), while using weight sharing significantly reduces the required number of samples to $1{,}000$ (blue and red solid lines). Notably, the initial test FER with weight sharing is already low even before training. A small amount of fine-tuning is enough to obtain the optimal weight set. As a result, the combination of transfer learning and weight sharing reduces the required number of samples from $50{,}000$ to $1{,}000$, thereby accelerating the sampling process by a factor of $50$.
Table VI presents a range of experiments to show the effectiveness of transfer learning across different code types, code rates, channels, and decoders. The target conditions for transfer learning are the WiMAX LDPC code with code rate of $3/4$, the AWGN channel, and the 5-bit quantized decoder, which have been assumed as a running example so far. Among them, we modify one condition at a time and derive the pre-trained weights. For the code type, we use the pre-trained weights from the WiFi LDPC code, whose results are also in Fig. 8. For all cases in Table VI, the fine tuning with $1{,}000$ samples improves the test FER compared to the baseline FER with merely transferring of weights, showing the effectiveness of transfer learning.
In addition, these results demonstrate the flexibility of the proposed boosting learning method, which is applicable to various code types, channels, and decoders (i.e., WiFi LDPC code, Rayleigh channel, low-precision decoder). In the remainder of the paper, we will further explore its applicability to 5G LDPC codes, short length LDPC codes, alternative quantization schemes, and the BP decoder. Our method can also be extended to broader classes, such as layered decoding [46] and optimized quantization schemes [50], which are left for future work.
III-F Data Augmentation
TABLE VII: Comparison between the plain sampling method and the accelerated sampling method using transfer learning and data augmentation
@@ III-F Data Augmentation
$\{\overline{w}_{v_{p}}^{(\ell)},w^{(\ell)}_{c_{p}\rightarrow v_{p}}\}_{\ell}$ Non-transfer $50000$ $2\times 10^{-5}$ N/A N/A $2.5\times 10^{9}$ $1$ 0.209
@@
In this subsection, we introduce the use of importance sampling for data augmentation to further speed up the sampling process. Importance sampling was used in estimating error floor performance of LDPC codes [7, 15]. Under the assumption of the binary phase shift keying (BPSK) modulation and all-zero codeword, stronger noise to variable nodes (VNs) within the trapping set is introduced by subtracting $\beta$ from their received values $y_{i}$. (i.e., $y_{i}=1-\beta+n_{\rm ch}$ for $i$ within the trapping set, where $n_{\rm ch}$ is the channel noise.) This approach allows for the rapid evaluation of relatively high FER, which is used to estimate the actual FER. Thus, we can estimate extremely low actual FER with a small number of decoding trials.
@@
The proposed data augmentation method also generates intensified noise at specific bit positions. As depicted in Fig. 7, the data augmentation block processes $T$ input UC vectors to produce $T\times D$ augmented UC vectors, where $T=100$ and $T\times D=1{,}000$. For each input UC vector, the base decoding is performed to identify the positions of error bits. These error bit positions, denoted as $\mathcal{E}$, are identified at the iteration when the number of UCNs is minimal. Subsequently, we accumulate $D$ UC vectors using the base decoder under a modified channel model where received vectors are generated from the AWGN channel as usual, and received values for VNs in $\mathcal{E}$ are shifted by $-\beta$. This shift increases the probability of decoding failures and enables the rapid collection of $D$ additional UC vectors that share characteristics with the input UC vector. Each UC vector entering the data augmentation block is augmented into $D$ UC vectors, thereby generating a total of $T\times D$ UC vectors from $T$ input UC vectors. Note that the error bit positions do not necessarily form a trapping set, as they are identified through the actual decoding process. Since we do not rely on algorithms to find trapping sets, data augmentation is feasible for long codes and is fast enough to collect augmented UC vectors.
@@
Table VII shows how transfer learning and data augmentation accelerates the sampling process. Without weight sharing, collecting $50{,}000$ UC vectors (indicated in ($a$)) with FER of the base decoder (indicated in FER ($b$)) requires $a/b=2.5\times 10^{9}$ decoding trials using the plain sampling method. By applying the dynamic weight sharing technique $\{\overline{w}^{(\ell)},w^{(\ell)},{\hat{w}}^{(\ell)}\}_{\ell}$, the required number of UC vectors is reduced to $10{,}000$ without transfer learning and further to $1{,}000$ with the addition of transfer learning. For data augmentation, Table VII presents results for $T=100$, $D=10$, and various $\beta$ values. With data augmentation, the augmented UC vectors ($c$) equal $T\times D=1{,}000$, and FER ($d$) represents FER under the modified channel with intensified noise. The decoding trials in data augmentation are $c/d$. Thus, the total number of decoding trials is given by $a/b+c/d$. The weight sharing technique alone reduces the required trials by a factor of $5$ and accelerates the sampling speed by the same factor. Transfer learning adds a 10-fold increase in speed, and this gain is further increased by data augmentation. As $\beta$ increases, FER ($d$) in data augmentation increases, giving more gains. However, augmented samples with high $\beta$ might differ in characteristics from those collected through plain sampling, which leads to a degraded test FER performance. Therefore, we opt for $\beta=0.7$, achieving a ×$\times 361$ increase in sampling speed without significant loss in the test FER. Boosting learning accompanied by transfer learning and data augmentation would be highly useful, especially in environments that need to support numerous code rates and lengths, such as 5G LDPC codes.
@@
Additionally, to eliminate the need for validation samples, we adopt the step decaying learning rate strategy. Until now, $5{,}000$ UC vectors have been allocated for validation samples. However, since the NMS decoder is not prone to overfitting on training data, the weights selected by validation samples have similar performance to those at the last epoch. To further enhance stability in convergence, we use the step decaying learning rate, starting from $0.001$ and halving it every $20$ epochs, and select the weights at the last epoch without validation samples.